# Enhancing Gravitational Wave Parameter Estimation with Non-Linear Memory: Breaking the Distance-Inclination Degeneracy

### v1

@@ IV.3 Case of LIGO ${}^{\sharp}$
Figure 8: Panel (a) shows Relative fractional improvement of MSE ($\Delta\mathrm{MSE}$) for full aligned-spin prior, with ${}^{\sharp}$ sensitivity curve. 15 injections are performed with 5 luminosity distances and 3 inclination angles (0, $\pi/4$, and $\pi/2$). The yellow lines are the SNR contour of the injected signal. Panel (b) and (c) show contour plots of select results with LIGO O5 and LIGO ${}^{\sharp}$ sensitivity curves at the same distance ($D_{L}=400/600$ Mpc) and inclination ($\iota=\pi/2$) in the same plot. The dashed lines indicate the injected values.
@@
From the exploration in the previous sections, we see that to break the degeneracy between inclination and luminosity distance, a high SNR greater than 60 is required. This means that for LIGO O5 sensitivities the distance should be closer than 300 - 600 Mpc, depending on the total mass and inclination angle. These kinds of close events are very few seen in O3 [1]. Therefore, we continue our exploration with the future upgrade of LIGO ${}^{\sharp}$ [43]. Using the sensitivity curve of ${}^{\sharp}$ is not simply equivalent to increasing the SNR of the overall signal, since the ${}^{\sharp}$ will also increase the sensitivity in the low frequency more, where most of the energy of memory lives.
@@
We performed similar injection sets for parameters as in the previous section. Due to the limited computational resources, we reduced the injection parameter space to 15 sets, with 3 inclination angles (0, $\pi/4$, and $\pi/2$) for 5 different distances. The injected distance started from 400 Mpc because the sensitivity of ${}^{\sharp}$ is much higher. The SNR below a distance of 400 Mpc is too high, which will cost too much time to finish.
@@
The $\Delta\mathrm{MSE}$ are plotted in Fig. 8a. This graph displays the improvement of MSE up to 90%. Even at around 1 Gpc, $\Delta\mathrm{MSE}$ can get around 10% - 20% of improvement. This means at that sensitivity, ignoring the non-linear memory effect will introduce significant systematics in the distance measurement for most of the significant events. This is estimated to affect nearly 30% of events reported in the GWTC-3 catalogue [1] assuming they were detected by ${}^{\sharp}$, taking the mean measured distance as the “true” distance to the source.
@@
We also plotted the comparison of inclination-distance posteriors for 2 select injection sets in both O5 and ${}^{\sharp}$ sensitivity curve in Fig. 8b and 8c. Fig. 8b shows the inclination-distance contour for the edge-on system at 600 Mpc. We can see that under ${}^{\sharp}$ sensitivity, the distance peaks at the injected value. Although the result using O5 sensitivity shows improvement in the MSE compared to the model without non-linear memory Fig. 7a, it is still largely unconstraint.
@@
Fig. 8c shows the inclination-distance contour for the edge-on system at 400 Mpc. For the O5 sensitivity, the SNR is around 50 at this distance. This comparable low SNR provides an improvement in the distance measurement seen from 7a, but still not enough to get rid of the bimodality. However, for ${}^{\sharp}$ the SNR is 90, and both inclination and distance are nicely peaked at the injected values.
@@
To conclude, for LIGO ${}^{\sharp}$, we will need memory to resolve distance-inclination degeneracy since we anticipate that the majority of events will have an SNR greater than 60, based on estimates using the GWTC-3 catalogue. Besides, for LIGO ${}^{\sharp}$, the $\rho_{20}^{\perp}>2$ when the distance is closer than 900 Mpc for total mass 40 $M_{\odot}$ binaries. For the same system at a distance of 600 Mpc, the $\rho_{20}^{\perp}$ is around 3. For now, events below a distance of 600 Mpc account for 12% of our total events in the GWTC-3 catalogue taking the mean measured distance, and those below 900 Mpc account for 24%.
@@ V Conclusions
In the presented study, we have conducted a detailed investigation into the utility of non-linear memory for improving parameter estimation within the context of gravitational wave astronomy, specifically focusing on its ability to resolve degeneracy in larger parameter spaces using Bayes inference injection methods. Initial assessments with a reduced prior setup were expanded to full-parameter recovery for aligned-spin systems, enabling a comprehensive evaluation of the impact of non-linear memory on enhancing the measurement of luminosity distance and inclination angle — key parameters for the accurate cosmological or astrophysical studies. The inclusion of non-linear memory has been shown to offer notable improvements in distance constraints, which is of significant importance for future observations with the upgraded LIGO ${}^{\sharp}$.
Our findings indicate that non-linear memory is particularly effective in breaking the degeneracy between inclination and luminosity distance in high signal-to-noise ratio (SNR $>$ 60) cases. The benefit of incorporating non-linear memory is more apparent at closer distances and for non-precessing and symmetric mass systems with higher SNRs. We also note that the (4,4) mode will have similar properties for the symmetric mass system but the inlication angle properties of the (4,4) mode are not as different from the (2,2) mode. The inclusion of (4,4) mode with (2,0) mode will overall improve the inclination angle measurement even for the intermediate values of inclinations as well which is not the case with just (2,0) mode as it improves mostly in the edge-on and face-on systems. The analysis suggests that neglecting non-linear memory effects in high SNR scenarios could lead to systematics in the estimated values of distance and inclination angles.
Further exploration into the implications of LIGO ${}^{\sharp}$’s upgraded sensitivity demonstrates the potential for enhanced parameter estimation accuracy, particularly due to its increased sensitivity at lower frequencies, where non-linear memory effects are more pronounced. The study suggests that with the ${}^{\sharp}$ sensitivity, omitting non-linear memory from the analysis could result in significant systematics in distance measurements for detectable events, even at distances up to about 1 Gpc. This is expected to affect almost 30% of the events in the GWTC-3 catalog based on mean measured distance assuming they were detected by ${}^{\sharp}$. This enhancement is attributed not only to the higher SNR provided by ${}^{\sharp}$ but also to its improved low-frequency sensitivity, crucial for capturing the entirety of non-linear memory effects.
@@
In summary, this research highlights the role of non-linear memory in refining the precision of gravitational wave parameter estimation. As we advance toward the LIGO ${}^{\sharp}$ phase and third-generation detectors, integrating non-linear memory into parameter estimation frameworks will be essential for the accurate localization and characterization of gravitational wave sources. This effort will contribute to the ongoing development of gravitational wave astronomy, enhancing our ability to analyze and interpret the signals from these cosmic phenomena.
@@ Acknowledgements.
We thank Shun Cheung and Paul Lasky for their constructive comments. Y. Xu is supported by the China Scholarship Council. S. Tiwari is supported by the Swiss National Science Foundation (SNSF) Ambizione Grant Number: PZ00P2-202204. C. Garciá-Quiroś is supported by the Swiss National Science Foundation (SNSF) grant Sinergia 213497. E. Hamilton was supported in part by the Swiss National Science Foundation (SNSF) grant IZCOZ0-189876 and by the UZH Postdoc Grant (Forschungskredit). Maria Rosselló-Sastre is supported by the Spanish Ministry of Universities via an FPU doctoral grant (FPU21/05009). E. Hamilton, S. Husa and M. Rosselló-Sastre were also supported in part by the Universitat de les Illes Balears (UIB); the Spanish Agencia Estatal de Investigación grants PID2022-138626NB-I00, PID2019-106416GB-I00, RED2022-134204-E, RED2022-134411-T, funded by MCIN/AEI/10.13039/501100011033; the MCIN with funding from the European Union NextGenerationEU/PRTR (PRTR-C17.I1); Comunitat Auton‘oma de les Illes Balears through the Direcció General de Recerca, Innovació I Transformació Digital with funds from the Tourist Stay Tax Law (PDR2020/11 - ITS2017-006), the Conselleria d’Economia, Hisenda i Innovació grant numbers SINCO2022/18146 and SINCO2022/6719, co-financed by the European Union and FEDER Operational Program 2021-2027 of the Balearic Islands; the “ERDF A way of making Europe”. E. Hamilton is supported in part by the Spanish Ministerio de Ciencia e Innovación (Fondos MRR) - Conselleria de Fons Europeus, Universitat i Cultura with funds from the European Union NextGenerationEU (PRTR-C17.I1) through the project ‘Tecnologías avanzadas para la exploración del universo y sus componentes’ (ref. SINCO2022/6719 ; PI: Alicia Sintes, University of the Balearic Islands, Spain). The authors are grateful for computational resources provided by Cardiff University supported by STFC grant ST/I006285/1, the LIGO Laboratory supported by National Science Foundation Grants PHY-0757058 and PHY-0823459, and ITP, Lanzhou University supported by National Key R&D Program of China grant No. 2021YFC2203003. This paper has document number LIGO-P2400063. This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.

### v2

@@ IV.3 Case of LIGO A
Figure 9: Panel (a) shows Relative fractional improvement of MSE ($\Delta\mathrm{MSE}$) for full aligned-spin prior, with A sensitivity curve. 15 injections are performed with 5 luminosity distances and 3 inclination angles (0, $\pi/4$, and $\pi/2$). The yellow lines are the SNR contour of the injected signal. Panel (b) and (c) show contour plots of select results with LIGO O5 and LIGO A sensitivity curves at the same distance ($D_{L}=400/600$ Mpc) and inclination ($\iota=\pi/2$) in the same plot. The dashed lines indicate the injected values.
@@
From the exploration in the previous sections, we see that to break the degeneracy between inclination and luminosity distance, a high SNR greater than 90 is required. This means that for LIGO O5 sensitivities the distance should be closer than 300 - 600 Mpc, depending on the total mass and inclination angle. These kinds of close events are very few seen in O3 [1]. Therefore, we continue our exploration with the future upgrade of LIGO A [52]. Using the sensitivity curve of A is not simply equivalent to increasing the SNR of the overall signal, since the A will also increase the sensitivity in the low frequency more, where most of the energy of memory lives.
@@
We performed similar injection sets for parameters as in the previous section. Due to the limited computational resources, we reduced the injection parameter space to 15 sets, with 3 inclination angles (0, $\pi/4$, and $\pi/2$) for 5 different distances. The injected distance started from 400 Mpc because the sensitivity of A is much higher. The SNR below a distance of 400 Mpc is too high, which will cost too much time to finish.
@@
The $\Delta\mathrm{MSE}$ are plotted in Fig. 9a. This graph displays the improvement of MSE up to 90%. Even at around 1 Gpc, $\Delta\mathrm{MSE}$ can get around 10% - 20% of improvement. This means at that sensitivity, ignoring the non-linear memory effect will introduce significant systematics in the distance measurement for most of the significant events. This is estimated to affect nearly 30% of events reported in the GWTC-3 catalogue [1] assuming they were detected by LIGO-A, taking the mean measured distance as the “true” distance to the source.
@@
We also plotted the comparison of inclination-distance posteriors for 2 select injection sets in both O5 and A sensitivity curve in Fig. 9b and 9c. Fig. 9b shows the inclination-distance contour for the edge-on system at 600 Mpc. We can see that under A sensitivity, the distance peaks at the injected value. Although the result using O5 sensitivity shows improvement in the MSE compared to the model without non-linear memory Fig. 8a, it is still largely unconstraint.
@@
Fig. 9c shows the inclination-distance contour for the edge-on system at 400 Mpc. For the O5 sensitivity, the SNR is around 80 at this distance. This comparable low SNR provides an improvement in the distance measurement seen from 8a, but still not enough to get rid of the bimodality. However, for A the SNR is 150, and both inclination and distance are nicely peaked at the injected values.
@@
To conclude, for LIGO A, we will need memory to resolve distance-inclination degeneracy since we anticipate that the majority of significant events will have a distance closer than 1 Gpc, based on estimates using the GWTC-3 catalogue. Besides, for LIGO A, the $\rho_{20}^{\perp}>2.7$ when the distance is closer than 1 Gpc for total mass 40 $M_{\odot}$ binaries. For the same system at a distance of 600 Mpc, the $\rho_{20}^{\perp}$ is around 4.5. For now, events below a distance of 600 Mpc account for 12% of our total events in the GWTC-3 catalogue taking the mean measured distance, and those below 1 Gpc account for 29%.
@@ V Conclusions
In the presented study, we have conducted a detailed investigation into the utility of non-linear memory in (2,0) mode contributed from dominant (2,2) oscillatory mode for improving parameter estimation within the context of gravitational wave astronomy, specifically focusing on its ability to resolve degeneracy in larger parameter spaces using Bayes inference injection methods. Initial assessments with a reduced prior setup were expanded to full-parameter recovery for aligned-spin systems, enabling a comprehensive evaluation of the impact of non-linear memory on enhancing the measurement of luminosity distance and inclination angle — key parameters for the accurate cosmological or astrophysical studies. The inclusion of non-linear memory has been shown to offer notable improvements in distance constraints, which is of significant importance for future observations with the upgraded LIGO A.
Our findings indicate that non-linear memory is particularly effective in breaking the degeneracy between inclination and luminosity distance in high signal-to-noise ratio (SNR $>$ 90) cases for LIGO-O5. The benefit of incorporating non-linear memory is more apparent at closer distances and for non-precessing and symmetric mass systems with higher SNRs. We also note that the (4,4) mode will have similar properties for the symmetric mass system but the inlication angle properties of the (4,4) mode are not as different from the (2,2) mode. The inclusion of (4,4) mode with (2,0) mode will overall improve the inclination angle measurement even for the intermediate values of inclinations as well which is not the case with just (2,0) mode as it improves mostly in the edge-on and face-on systems. The analysis suggests that neglecting non-linear memory effects in high SNR scenarios could lead to systematics in the estimated values of distance and inclination angles.
Further exploration into the implications of LIGO A’s upgraded sensitivity demonstrates the potential for enhanced parameter estimation accuracy, particularly due to its increased sensitivity at lower frequencies, where non-linear memory effects are more pronounced. The study suggests that with the A sensitivity, omitting non-linear memory from the analysis could result in significant systematics in distance measurements for detectable events, even at distances up to about 1 Gpc. This is expected to affect almost 30% of the events in the GWTC-3 catalog based on mean measured distance assuming they were detected by LIGO-A. This enhancement is attributed not only to the higher SNR provided by A but also to its improved low-frequency sensitivity, crucial for capturing the entirety of non-linear memory effects.
@@
In summary, this research highlights the role of non-linear memory in refining the precision of gravitational wave parameter estimation. As we advance toward the LIGO A phase and third-generation detectors, integrating non-linear memory into parameter estimation frameworks will be essential for the accurate localization and characterization of gravitational wave sources. This effort will contribute to the ongoing development of gravitational wave astronomy, enhancing our ability to analyze and interpret the signals from these cosmic phenomena. Additionally, a newly developed phenomenological waveform model for the memory and oscillatory part in the (2,0) mode will facilitate the routine inclusion of memory effects in the future data analysis [54].
@@ Acknowledgements.
We thank Shun Cheung, Paul Lasky, and the referee for their constructive comments. Y. Xu is supported by the China Scholarship Council. S. Tiwari is supported by the Swiss National Science Foundation (SNSF) Ambizione Grant Number: PZ00P2-202204. C. Garciá-Quiroś is supported by the Swiss National Science Foundation (SNSF) grant Sinergia 213497. E. Hamilton was supported in part by the Swiss National Science Foundation (SNSF) grant IZCOZ0-189876 and by the UZH Postdoc Grant (Forschungskredit). Maria Rosselló-Sastre is supported by the Spanish Ministry of Universities via an FPU doctoral grant (FPU21/05009). E. Hamilton, S. Husa and M. Rosselló-Sastre were also supported in part by the Universitat de les Illes Balears (UIB); the Spanish Agencia Estatal de Investigación grants PID2022-138626NB-I00, PID2019-106416GB-I00, RED2022-134204-E, RED2022-134411-T, funded by MCIN/AEI/10.13039/501100011033; the MCIN with funding from the European Union NextGenerationEU/PRTR (PRTR-C17.I1); Comunitat Autonòma de les Illes Balears through the Direcció General de Recerca, Innovació I Transformació Digital with funds from the Tourist Stay Tax Law (PDR2020/11 - ITS2017-006), the Conselleria d’Economia, Hisenda i Innovació grant numbers SINCO2022/18146 and SINCO2022/6719, co-financed by the European Union and FEDER Operational Program 2021-2027 of the Balearic Islands; the “ERDF A way of making Europe”. E. Hamilton is supported in part by the Spanish Ministerio de Ciencia e Innovación (Fondos MRR) - Conselleria de Fons Europeus, Universitat i Cultura with funds from the European Union NextGenerationEU (PRTR-C17.I1) through the project ‘Tecnologías avanzadas para la exploración del universo y sus componentes’ (ref. SINCO2022/6719 ; PI: Alicia Sintes, University of the Balearic Islands, Spain). The authors are grateful for computational resources provided by Cardiff University supported by STFC grant ST/I006285/1, the LIGO Laboratory supported by National Science Foundation Grants PHY-0757058 and PHY-0823459, and ITP, Lanzhou University supported by National Key R&D Program of China grant No. 2021YFC2203003. This paper has document number LIGO-P2400063. This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.
Appendix A The effect of aligned-spin magnitude
The amplitude of aligned spin influences the memory amplitude, potentially affecting parameter estimation with memory. In this section, we conducted a series of injections similar to those described in Section IV.1. We injected 25 sets of parameters, comprising 5 values of $s_{1z}$ (evenly spaced from 0 to 0.8) and 5 inclination angles (evenly spaced from 0 to $\pi/2$). The other parameters remain the same as Section IV.1 and the aligned spin magnitudes $a_{1}$ and $a_{2}$ are varied in the prior. Figure 11 illustrates the improvement in distance measurement $\Delta\mathrm{MSE}$. The results indicate that for $s_{1z}$ values less than 0.6, the impact on distance measurement is relatively modest with (2,2) mode and its memory contribution in (2,0). However, a substantial improvement in the $\Delta\mathrm{MSE}$ is observed when the spin reaches 0.8. The decrease in $\Delta\mathrm{MSE}$ around $\iota=\pi/4$ is due to the smaller difference compared to the face-on and edge-on cases, with respect to the less informative IMRPhenomT recovery as shown in the upper panel of Fig. 6.
Figure 11: The plot shows the relative fraction improvement $\Delta\mathrm{MSE}$ for reduce prior, varying the spin magnitude of $s_{1z}$. Twenty-five injections are performed with five inclination angles and five $s_{1z}$ values.